OpenStudy (anonymous):

find (fog)(x) if f(x)=x-x^2 and g(x)=3x+2

OpenStudy (anonymous):

first get rid of that circle and write \[f\circ g(x)=f(g(x))\] then work from the inside out replace \(g(x)\) by \(3x+2\) and write \[f(g(x))=f(3x+2)\]

OpenStudy (anonymous):

then, wherever you see an \(x\) in \(f(x)\) replace it by \(3x+2\) to get \[f(3x+2)=3x+2-(3x+2)^2\]